OpenStudy (ac3):

Find all solutions of cosh(z) = 0

OpenStudy (ac3):

I got z = ln(+or- 1) + ipi/2 + i2kpi, back of the book says it's (1+or-2k)ipi. Are the answers equivalent?

OpenStudy (ac3):

Just went through it again, now I got z = ln(i) + 2ikpi, k = 0,1,2.....

OpenStudy (irishboy123):

you could have a go with: \(\cosh z = \cos iz \)